OpenStudy (anonymous):

solve 2x+2(x+3)=2(x+7) but i need to know the steps it helps me out...plz

OpenStudy (anonymous):

first distribute 2x+(2x+6)=2x+14 then simplify 4x +6= 2x+14 simplify more 2x+6=14 more 2x=14-6=8 x=4

OpenStudy (ksaimouli):

x=4

OpenStudy (anonymous):

2x+2x+6=2x+14 Distributive property 4x+6=2x+14 Addition of like terms on the same side of = -2x -2x Subtraction prop of equality 2x+6=14 -6 -6 ditto 2x=8 2x/2=8/2 Division prop of equality x=4